### v1

# Scalable Vision Language Model Training via High Quality Data Curation

Abstract
In this paper, we introduce SAIL-VL (ScAlable Vision Language Model TraIning via High QuaLity Data Curation), an open-source vision language model (VLM) of state-of-the-art (SOTA) performance with 2B parameters. We introduce three key improvements that contribute to SAIL-VL’s leading performance: (1) Scalable high-quality visual understanding data construction: We implement a visual understanding data construction pipeline, which enables hundred-million-scale high-quality recaption data annotation. Equipped with this pipeline, we curate SAIL-Caption, a large-scale caption dataset with large quantity and the highest data quality compared with opensource caption datasets. (2) Scalable Pretraining with High-Quality Visual Understanding Data: We scale SAIL-VL’s pretraining budget up to 131B tokens and show that even a 2B VLM benefits from scaled up training data sizes, exhibiting expected data size scaling laws in visual understanding and instruction following performance. (3) Scalable SFT via quantity and quality scaling: We introduce general guidance for instruction data curation to scale up instruction data continuously, allowing us to construct a large SFT dataset with the highest quality. To further improve SAIL-VL’s performance, we propose quality scaling, a multi-stage training recipe with curriculum learning, to improve model performance scaling curves w.r.t. data sizes from logarithmic to be near-linear.
SAIL-VL obtains the highest average score in 19 commonly used benchmarks in our evaluation and achieves top1 performance among VLMs of comparable sizes on OpenCompass (https://rank.opencompass.org.cn/leaderboard-multimodal). We release our SAIL-VL-2B model at HuggingFace (https://huggingface.co/ BytedanceDouyinContent/SAIL-VL-2B).
Scalable Vision Language Model Training via High Quality Data Curation
Hongyuan Dong*, Zijian Kang*, Weijie Yin*, Xiao Liang, Chao Feng, Jiao Ran ByteDance Douyin Content Group {donghongyuan.dousia, zijian.kang, yinweijie} {liangxiao.ilx, chaofeng.zz, ranjiao}@bytedance.com
footnotetext: * Equal contribution.footnotetext: †Email corresponding
1 Introduction
Researches in vision large language models (VLMs) Liu et al. (2024a); Li et al. (2024a); Yao et al. (2024); Wang et al. (2024); Gu et al. (2024); Chen et al. (2024d, c) have made significant progress in recent years, facilitating various vision tasks via language interactions. Due to the memory and computational constraints in model deployment, training compact VLMs with robust visual comprehension performance has become a popular research field recently Marafioti et al. (2024); Chen et al. (2023a); Yao et al. (2024); Li et al. (2024c); Gao et al. (2024). However, how to make optimal use of publicly available resources to unlock the potential of compact VLMs remains an unanswered question. We attribute the suboptimal performance of recent lightweight vision language models to two aspects: limited fundamental visual understanding abilities and unsatisfactory visual instruction following performance.
The fundamental visual understanding abilities of VLMs are typically established via large-scale pretraining. VLM pretraining necessitates not only substantial training budgets, but also a sufficient amount of high-quality visual understanding data to take effect. Recently proposed VLMs, such as LLaVA series Liu et al. (2024a); Li et al. (2024a); Chen et al. (2024a), conduct light-weight pretraining with a limited amount of low-quality caption data, and therefore suffer from suboptimal visual understanding abilities which hinder subsequent visual instruction tuning. MiniCPM-V-2.5 Yao et al. (2024) and Qwen2-VL Wang et al. (2024) allocate hundreds of billions of tokens’ computation budgets to the pretraining stage, but the limited visual understanding data quality hinders these models from achieving stronger visual understanding capabilities. More importantly, despite the large amount of resources consumed in VLM pretraining, existing works do not provide reliable conclusions to understand how pretraining budgets and data quality influence VLM performance.
During supervised fine-tuning (SFT) stage, VLM’s visual understanding capabilities are generalized to instruction following tasks. However, how to make optimal use of high-quality visual instruction tuning datasets remains unexplored. Recent works primarily focus on adjusting the data distribution across various domains and formats Li et al. (2024a); Chen et al. (2024d, c); Yao et al. (2024). Infinity-MM Gu et al. (2024) proposes multi-stage SFT to improve model performance, and obtain promising performance scaling results. Nevertheless, there still lacks widely acknowledged guidelines to determine the distribution of SFT datasets and training stages.
To address the above issues, we propose SAIL-VL, an opensource vision language model with state-of-the-art (SOTA) visual understanding and instruction following performance. SAIL-VL is trained through five stages: pretrain-alignment, pretrain-advance, SFT-knowledge, SFT-instruction and SFT-preference. In the first two pretraining stages, we establish SAIL-VL’s fundamental visual understanding abilities with large-scale pretraining. To explore how pretraining computation budgets and data quality influence VLM’s visual understanding abilities, we scale up VLM pretraining to 131B tokens with SAIL-Caption, our synthesized visual understanding dataset, which is of large magnitude and the highest quality compared to other publicly available caption datasets. During the following three SFT stages, we train SAIL-VL through three separate stages with different training objectives. Moreover, we elaborate guidelines to select SFT datasets with balanced distribution and the guidelines for the design of VLM SFT stages to further improve model performance. Trained through five stages, our SAIL-VL-2B achieves top-1 performance on OpenCompass leaderboard compared to VLMs of similar sizes, and obtains the highest average score in 19 commonly used benchmarks in our evaluation.
We summarize the key contribution of this research as below:
(1) We implement a data construction pipeline for scalable high-quality visual understanding data construction. Equipped with this pipeline, we construct SAIL-Caption dataset, which is of large magnitude and the highest quality compared with opensource caption datasets.
(2) We scale up SAIL-VL’s pretraining data size to 131B tokens, and report logarithmic model performance scaling laws w.r.t. training data size and training budgets. To the best of our knowledge, this is the first time that data size scaling laws for VLM pretraining are proposed and discussed.
(3) We elaborate the guidelines for the design of VLM SFT strategies and the distribution of visual instruction tuning datasets, which enable SAIL-VL-2B to achieve top1 performance on OpenCompass leaderboard and obtain the highest average score in 19 commonly used VLM benchmarks compared with previous SOTA models of comparable sizes.
2 Model Training Pipeline
In this section, we introduce SAIL-VL’s training strategy shown in Fig 1. Starting from InternViT-300M Chen et al. (2023b) and Qwen-2.5-1.5B Team (2024b), SAIL-VL is pretrained for visual understanding in the first two pretraining stages and then adapted to visual instruction tasks in instruction tuning stages.
Figure 1: SAIL-VL’s overall model training pipeline.
2.1 Pretrain
During the pretraining stages, we gradually open model parameters for training to develop SAIL-VL’s fundamental visual understanding abilities from large-scale pretraining data.
Pretrain-Alignment.
We start from a randomly initialized multi-layer perceptron (MLP) module which bridges the representation spaces of visual features and language models. We train the vision-to-language projector with approximately 16B tokens in this stage, equipping SAIL-VL with basic visual understanding abilities. Detail caption data and OCR data are used to train the model in this stage.
Pretrain-Advance.
After the alignment pretrain stage, we unlock the visual encoder of SAIL-VL and train the model with approximately 131B tokens. We include more OCR data and a larger proportion of SAIL-Caption for model training. We do not use the entire SAIL-Caption dataset but a subset of it, selected to maintain an even distribution across different data sources, thereby ensuring diversity in data distribution.
2.2 Instruction Tuning Stage (SFT)
We train all parameters of SAIL-VL through three SFT stages in a curriculum learning fashion with progressively higher-quality training data.
SFT-Knowledge.
In this stage, SAIL-VL learns basic instruction-following abilities and ingests world knowledge from the training data. We use the data from Infinity-MM Stage2 Gu et al. (2024) for SFT-Knowledge stage training, which is of the largest quantity and medium quality among training datasets for the three SFT stages.
SFT-Instruction.
We further optimize the SAIL-VL model to obtain enhanced visual instruction following capabilities with a balanced distribution across downstream domains. We use a 12M customized visual instruction tuning dataset, including more open-ended question-answering, complex OCR VQA, and other visual instruction following data with more diverse distribution, prioritizing data quality over sheer volume.
SFT-Preference.
During the last stage, we train SAIL-VL on scarce yet high-quality data with open-ended answers aligned with human preference, including LLaVA Li et al. (2024a) SFT data, Molmo Caption Deitke et al. (2024), and Infinity-MM Stage4 Gu et al. (2024). Preference learning data for this stage is more complex and harder to collect, and is therefore used in the last SFT stage for training effectiveness.
3 Towards Scalable VLM Training
Dataset Language # Sample Avg. Len. Uni. 1-gram Uni. 2-gram Uni. 3-gram Uni. Noun Uni. Verb Uni. Adj.
[…]
DataComp-LLaVA-Caption (2024b) EN 940,891,257 48.37 1.661 8.400 18.23 0.4728 0.0333 0.2801
[…]
SA1B-QwenVL-Caption (2024b) CN 8,631,495 130.3 0.2819 7.450 22.08 0.5797 0.1828 0.0378
[…]
Table 1: Statistics of SAIL-Caption and other opensource caption datasets. We employ NLTK Bird (2006) and Jieba Sun to perform text segment and part-of-speech tagging for English and Chinese captions, respectively. “Avg. Len.” stands for “average length” and “Uni.” denotes “unique”.
We elaborate the design of our scalable model training pipeline: We first establish a scalable high-quality data construction pipeline, and then validate the data scaling laws observed in pretraining with this data. Additionally, we introduce the scaling law for instruction tuning, a data curation pipeline for quantity scaling, and a multi-stage training pipeline for quality scaling.
3.1 Scalable High-Quality Visual Understanding Data Construction
We illustrate our scalable high-quality detail caption data construction pipeline in Figure 1, which can be described as the following four steps.
Data collection.
We collect source data from a wide range of publicly available datasets to ensure our visual understanding datasets possess sufficiently diverse distribution. Our source datasets include LAION-COCO Schuhmann et al. (2022), TextCaps Sidorov et al. (2020), SA1B Kirillov et al. (2023), and several other large-scale datasets.
Reference data curation.
We curate a small amount of reference data to train an efficient VLM to annotate high-quality visual understanding data efficiently at scale. We first select a subset of our collected source images with a balanced distribution across different source datasets, and then task powerful VLM APIs to annotate detail captions. It is worth noticing that we follow previous works Yu et al. (2024); Hong et al. (2024) to provide alt-text for reference data generation if possible, providing supplementary world knowledge for enhanced reference data quality.
Captioner model training.
Equipped with the high-quality reference data, we train a compact VLM for scalable caption data generation. We train SAIL-Captioner on the reference dataset, with alt-text optionally included in the prompt, enabling SAIL-Captioner to perform both captioning and recaptioning.
Scalable high-quality data construction.
In the final stage, we use SAIL-Captioner to annotate large-scale visual understanding datasets for model pretraining. We implement a data construction pipeline, which supports multi-task, multi-node, and multi-processing asynchronous annotation, to enable flexible data annotation resource allocation. LMDeploy Contributors (2023) is incorporated in our annotation system to further speed up the generation process.
SAIL-Caption.
[…]
To quantitatively demonstrate the advantages of the SAIL-Caption dataset over other open-source datasets, we curated a 500-case subset from each dataset for data quality evaluation. The detailed evaluation process is elaborated in Appendix B.1. Results show that SAIL-Caption-DataComp and SAIL-Caption-SA1B outperform previous baseline datasets by $20.35\%$ and $11.91\%$, respectively, exhibiting considerable data quality improvements. We also show visualized showcases of SAIL-Caption in Appendix B.2.
Figure 2: Scaling curves of SAIL-VL’s performance dynamics in pretrain-advance stage. We show model performance on all understanding benchmarks, caption tasks and OCR tasks, respectively.
Figure 3: Scaling curves of SAIL-VL’s performance dynamics in pretrain-advance stage. We show average model performance on all VLM benchmarks used in our experiments, general VQA tasks, OCR VQA tasks, respectively. The two sub-domains cover visual instruction following tasks for natural scenes and complicated documents, respectively. The detailed benchmark split for different domains is shown in Table 2.
3.2 Scalable VLM Pretraining With High-Quality Visual Understanding Data
In this part, we introduce the scaling laws observed in SAIL-VL large-scale pretraining. We scale up data size used in SAIL-VL pretrain-advance stage, with training budgets configured accordingly. For model checkpoints obtained at different pretraining steps, we conduct a lightweight annealing training to let model parameters converge to local optimal points, ensuring the stability of evaluation results.
3.2.1 Improving VLM Visual Understanding Performance via Data Size Scaling
SAIL-VL is trained through 16B tokens for initial vision-language alignment, and then ingests 131B tokens during the pretrain-advance stage, during which we investigate model performance dynamics. To evaluate the visual understanding performance of SAIL-VL, we establish an evaluation suite which covers fundamental visual understanding tasks such as detail caption generation Dong et al. (2024) and OCR detection Biten et al. (2022); Wang et al. (2020); Gupta et al. (2016); Kim et al. (2022). Details can be found in Appendix D.
We select checkpoints trained with exponentially larger data sizes from the pretrain-advance stage to inspect model performance dynamics. As shown in Figure 2, SAIL-VL’s overall performance, captioning performance, and OCR detection performance improve smoothly as the training proceeds. As pretraining data sizes and training budgets scale up exponentially, SAIL-VL’s visual understanding performance improves linearly. This observation unveils a promising prospect to scale up VLM pretraining data sizes for improved model performance.
3.2.2 Generalizing Visual Understanding Abilities to Instruction Following Tasks
To further investigate the effectiveness of SAIL-VL’s large-scale pretraining, we conduct visual instruction tuning with model checkpoints collected from the pretrain-advance stage. Specifically, we curate visual instruction tuning data from the SFT-instruction stage, and evaluate the resulting model checkpoints on 19 commonly used VLM benchmarks to validate the scaling results.
As shown in Figure 3, SAIL-VL’s overall performance improves steadily as the data size and computation budget increase, and the dynamics can be plotted as a near-linear curve on an exponential horizontal axis. Even at the data point derived from the model checkpoint trained with 131B tokens, the scaling curve of SAIL-VL’s performance does not converge. We also show the scaling results on two widely used benchmark series: General VQA and OCR VQA. As the training budget scales up exponentially, SAIL-VL’s consistently develops more robust abilities to recognize visual elements and tackle instruction following tasks, leading to near-linear performance scaling curves.
3.3 Scaling up VLM Instruction Tuning
The instruction tuning (SFT) stage is crucial for VLMs’ downstream performance. While the visual instruction tuning data from the internet is abundant, high-quality data is still scarce. To solve this problem, we introduce several guidelines for visual instruction tuning data curation in Section 3.3.1, and demonstrate model performance scaling laws of the curriculum SFT strategy with progressively improved data quality in Section 3.3.2.
3.3.1 Instruction Data Curation for Quantity Scaling
In this section, we elaborate on the guidelines for visual instruction tuning data curation and demonstrate their effectiveness by training with our customized SFT data.
Efficient data quality evaluation.
Firstly, we propose an intuitive way for data quality evaluation. For any dataset, we train a model on a small subset of data with a fixed quantity (2M samples) for efficiency, and then conduct holistic evaluation on a series of benchmarks. The evaluation results reflect training data quality comprehensively. The SFT data distribution is decided with reference to the data quality evaluation results.
Detecting low-quality components from SFT data mixing.
To detect which part of visual instruction data undermines model performance the most, we propose Composition Evaluation, which identifies data quality of SFT data components from model performance variations by halving each part of the data. We conduct composition evaluation on each SFT data category, and repeat this process until no further improvement is observed. Equipped with the proposed composition evaluation strategy, we improve SAIL-VL’s SFT-instruction training data distribution across existing SFT data sources to achieve improved model performance.
Figure 4: Scaling curves of model performance trained on our SAIL-Instruct dataset, LLaVA-OneVision Li et al. (2024a) single image SFT data, and datasets from Infinity-MM Gu et al. (2024).
Identifying incoming visual instruction tuning data quality.
For incoming visual instruction tuning data to be incorporated, we explore scaling up SFT data sizes without compromising overall data quality. We propose the Incremental Evaluation strategy, which identifies the quality of an incoming visual instruction tuning dataset from model performance variations with the incoming dataset included in the training data. This simple-yet-effective approach enables us to continuously expand data quantity while preserving its quality.
Figure 5: Model performance dynamics of the quality scaling and all-in-one (AIO) training strategy. “AIO learning” incorporates all three-stage SFT data into a single training loop.
Quantity Scaling.
Based on this methodology, we collect a large amount of high-quality visual instruction tuning data. As shown in Figure 4, model performance scales up stably as the data quantity grows from 2M to 12M, depicting a logarithmic scaling curve as training data quantity increases. Training with datasets of different quality results in varying model performance. Among these datasets, our SAIL-Instruct dataset achieves the best quality evaluation results, surpassing recently proposed alternatives either at the beginning of model training or at the end of the scaling curves.
3.3.2 Multi-stage Instruction Tuning with Quality Scaling
In this part, we introduce quality scaling, a curriculum learning strategy for VLM SFT, which starts with abundant low-quality data and switches to scarce high-quality data at the end of the training, to achieve even higher model performance than one-stage SFT.
Curriculum SFT with progressively improving data quality.
We train SAIL-VL through three SFT stages in a curriculum learning paradigm. SAIL-VL ingests world knowledge and develops basic visual instruction following abilities during the first two stages, and learns to tackle complex tasks in the last training stage. As elaborated in Section 2.2, visual instruction tuning data used in the three SFT stages differs in the following ways: (1) Datasets with more scarce data quantity are used for later stages. (2) Data formats become increasingly complex across the stages. (3) Data quality progressively improves through the training process. The three-stage model training strategy ensures effective utilization of scarce-yet-high-quality datasets, and therefore enables SAIL-VL to develop robust and comprehensive visual instruction following abilities.
Quality Scaling.
We show model performance dynamics derived from the three SFT stages in comparison with an all-in-one (AIO) training strategy in Figure 5. The model trained with the curriculum SFT strategy exhibits a near-linear performance scaling curve across training stages with progressively improved data quality, yielding better final model performance over AIO training. Model performance dynamics during AIO training can be described as a logarithmic curve with suboptimal performance. We attribute this performance inferiority to the inefficient data allocation strategy compared to our curriculum SFT approach.
4 Experiments
Benchmark SAIL-VL Qwen2-VL InternVL2 InternVL2.5 Aquila DeepSeekVL-2-Tiny
Overall Performance
Opensource Average 68.6 64.0 62.3 66.1 65.9 66.8
   General VQA 60.2 58.2 55.3 62.1 59.2 59.9
   OCR VQA 75.7 72.5 68.2 68.7 71.7 74.4
   Math&Knowledge 79.6 59.0 70.6 73.0 75.4 71.3
   Hallucination 66.9 62.9 61.7 66.2 62.9 63.6
OpenCompass Average 61.3 55.2 54.1 59.6 58.7 57.3
General VQA
MMStar (2024b) 56.6 46.3 50.5 53.5 54.7 49.9
MMBench (2024b) 74.4 68.8 70.3 73.1 74.4 68.3
MMMU (2024b) 42.0 39.9 34.2 40.7 44.1 39.6
MME (2024b) 2001 1923 1860 2091 1808 1911
SEEDBench (2024b) 74.7 72.0 70.9 73.4 73.9 72.5
RealWorldQA (2024b) 62.7 60.9 56.7 60.9 64.1 64.8
MMVet (2023) 47.0 51.2 40.4 61.7 42.7 52.8
LLaVABench (2024b) 52.3 57.5 52.8 59.0 54.8 62.8
OCR VQA
AI2D (2024b) 77.3 72.3 74.2 75.0 75.0 74.6
DocVQA (2024b) 88.3 88.7 86.0 87.4 85.0 88.6
InfoVQA (2024b) 66.5 63.4 57.5 61.6 60.5 63.8
ChartQA (2024b) 79.6 70.6 71.7 73.3 76.6 81.2
TextVQA (2024b) 77.3 78.8 73.4 76.4 76.4 80.5
OCRVQA (2024b) 57.2 54.3 36.2 28.3 51.3 51.4
OCRBench (2024b) 835 794 786 789 772 808
Math&Knowledge
MathVista (2024b) 63.4 45.0 46.8 51.1 59.4 54.5
ScienceQA (2024b) 95.8 73.0 94.4 94.9 91.4 88.1
Hallucination
HallusionBench (2024b) 46.0 38.3 38.2 42.5 42.1 38.4
POPE (2024b) 87.8 87.6 85.3 89.9 83.6 88.8
Table 2: Evaluation results of SAIL-VL and other opensource VLM with comparable sizes. “Opencompass Average” indicates average model performance on OpenCompass benchmarks, while “Opensource average” includes all opensource benchmarks listed in the table. Bold numbers indicate the best performance on the benchmark, while underlined ones are those ranked as the second.
4.1 Experiment Setup
Training framework.
We use PyTorch Paszke et al. (2019); Ansel et al. (2024) version 2.1.0 with CUDA NVIDIA (2024) 12.1 for model training. Deepspeed Rasley et al. (2020) version 0.14.5 is used for SAIL-VL training. Flash-attention Dao et al. (2022); Dao (2023) implemented for 910B NPU Huawei (2023) is leveraged for fast attention computation.
We process training data sequences with a stream accumulator, which packs sequences in a micro batch into a long sequence for model training. This strategy speeds up SAIL-VL model training by approximately $40\%$.
Training resources.
We conduct experiments with Huawei 910B x86 NPU. A total of 26,656 NPU hours are consumed to train SAIL-VL-2B, with pretraining stages and SFT stages consuming 15,564 and 10,992 NPU hours, respectively.
Baselines.
We compare our SAIL-VL-2B model with previous SOTA VLM baselines of comparable sizes, including Qwen2-VL-2B Wang et al. (2024), InternVL2-2B Team (2024a), InternVL2.5-2B Chen et al. (2024c), Aquila-2B Gu et al. (2024), and DeepSeekVL-2-2BWu et al. (2024).
Evaluation.
We evaluate SAIL-VL and baseline VLMs on a series of widely used benchmarks. We categorize the opensource benchmarks used in our evaluation pipeline into four types: General VQA, OCR VQA, Math&Knowledge, and Hallucination, covering VQA tasks on natural images/videos, OCR-related documents, as well as those involving complicated reasoning and world knowledge to tackle. All results are evaluated in our customized version of VLMEvalKit Duan et al. (2024), which may differ slightly from model performance reported in other papers or the official OpenCompass leaderboard.
4.2 Benchmark Results
SAIL-VL ourperforms previous SOTA VLMs with comparable sizes significantly.
We list the performance of SAIL-VL along with other opensource VLMs in Table 2. As the results show, SAIL-VL outperforms previous SOTA VLMs by a large margin w.r.t. average scores across all benchmarks or the OpenCompass benchmark collection, indicating robust visual understanding and instruction following abilities. SAIL-VL achieves the top-2 performance on 13 out of 19 benchmarks in our evaluation. Compared to previous SOTA VLMs of comparable sizes, SAIL-VL scores $1.8$ ($2.69\%\uparrow$) higher opensource average performance than DeepSeekVL-2-Tiny, and $1.7$ ($2.85\%\uparrow$) higher OpenCompass average score than InternVL-2.5, respectively.
SAIl-VL scores the highest in visual comprehension across in almost every domain.
It is worth noticing that SAIL-VL achieves leading performance across benchmarks of various domains. SAIL-VL exceeds previous SOTA VLMs by $1.3$ ($1.75\%$) in OCR VQA, $4.2$ ($5.57\%$) in math&knowledge tasks, and $0.7$ ($1.06\%$) in Hallucination benchmarks, respectively. SAIL-VL’s performance decays slightly in general VQA tasks, especially in the MMVet benchmark. We attribute this performance drop to the open-ended answering paradigm, which may introduce instability in evaluation because of the varying answering lengths of different VLMs. These results indicate that SAIL-VL develops powerful visual comprehension and language instruction abilities within its limited model size, demonstrating promising prospects in downstream applications.
5 Analysis
5.1 Training Pipeline Ablation
PT-adv SFT-knw SFT-inst SFT-prf OS Avg
68.6
67.7
67.7
66.2
67.1
Table 3: SAIL-VL’s performance with each training stage removed. ✓and ✗ indicate that this stage is retained and removed, respectively. “OS Avg” stands for average model performance on the 19 opensource benchmarks.
We conduct ablation experiments to show the necessity and effectiveness of our five-stage training pipeline. We experiment with each training stage removed except the pretrain-alignment stage, because a randomly initialized projector may cause ViT parameter distribution collapse during the subsequent training stages. Experiment results are shown in Table 3, where a consistent performance drop is witnessed with any training stage removed, validating the necessity and effectiveness of our training pipeline design.
5.2 Pretrain Data Quality Determines Pretrained Model Performance
We explore pretraining SAIL-VL with caption data of varying data quality. Specifically, we conduct lightweight experiments in the pretrain-advance stage, starting from the model checkpoint after the same alignment pretraining. A total of 16B tokens are consumed for these ablation studies, consisting of a fixed amount of OCR and caption data from different data sources. We evaluate model performance with our understanding benchmarks.
Caption Data Understanding Avg. Caption Avg. OCR Avg.
SAIL-Caption 52.37 51.19 53.54
SA1B-QwenVL-Caption 50.97 48.20 53.74
DataComp-LLaVA-Caption 48.90 44.92 52.88
Table 4: Visual understanding performance of model checkpoints pretrained with different data sources. We report models’ average visual understanding performance on caption and OCR benchmarks, respectively.
As shown in Table 4, the model trained with SAIL-Caption achieves significantly higher performance than those trained on other opensource caption datasets. The model performance evaluation results coincide smoothly with our data quality assessment findings: the higher the pretrain caption data quality, the better the pretrained model performs. Details about our data quality evaluation results are shown in Appendix 6.
5.3 Generalizing Visual Understanding Abilities to Instruction Following Tasks
Figure 6: The correlation between SAIL-VL pretrained checkpoints’ understanding performance and their performance on opensource benchmarks after SFT. “OS BMK Score” stands for average score on opensource benchmarks.
In this part, we investigate the correlation between VLM’s visual understanding and instruction following abilities. We train SAIL-VL pretrained checkpoints from the pretrain-advance stage with exponentially larger training data sizes, and train them through either LLaVA-Next Liu et al. (2024a) SFT data or our SFT-instruction data. We plot the correlation of pretrained models’ visual understanding performance and SFT models’ opensource benchmark performance in Figure 6. A notable correlation is observed across different training strategies. As the VLM gains stronger visual understanding abilities during pretraining, its visual instruction following performance after SFT is improved accordingly, even if trained with different visual instruction tuning datasets. We quantify this correlation with Pearson correlation ($\rho$) and coefficient of determination ($R^{2}$). It turns out that SAIL-VL’s pretrained visual understanding performance and SFT visual instruction following performance share a significant correlation with $\rho=0.91$ and $R^{2}=0.83$, validating the generalization of model abilities across training stages.
5.4 Data Composition and Quality Evaluation for Visual Instruction Tuning
Data Composition Quality Evaluation
Stages Size General OCR Math. Text Overall General OCR Math. Hall.
Knowledge Learning 21M 36% 43% 7% 14% 0.548 0.511 0.582 0.556 0.568
Instruction Learning 12M 44% 31% 9% 16% 0.609 0.542 0.657 0.681 0.631
Preference Learning 3.5M 76% 11% 10% 3% 0.606 0.560 0.658 0.595 0.613
Table 5: Composition and quality evaluation of dataset from each stage. Note that Math. for data composition includes data with math and reasoning, Math. for evaluation denotes Math &Knowledge, Hall. denotes Hallucination.
To demonstrate the differences between SFT stages, we show the data distribution of each stage along with the performance of models trained with each part of SFT data. As illustrated in Table 5, the composition of knowledge learning and instruction learning data is similar, but the resulting model performance improves significantly. The composition of preference learning data differs from data used in other stages, containing much more open question-answering samples. Although training with preference learning data does not yield the highest performance scores, model performance in general VQA tasks. With a small amount of high-quality OCR samples, the OCR evaluation score for preference learning is also high.
6 Related Works
6.1 Visual Understanding Data
Visual understanding data consists of vision modality contents and corresponding language depictions, and is regarded as the keystone to various vision and language model applications. Whether it is representation learning models like CLIP and its derivatives Radford et al. (2021); Jia et al. (2021); Shen et al. (2022); Cherti et al. (2023); Fang et al. (2023), generative models Wang et al. (2022); Li et al. (2021); Bao et al. (2022); Yu et al. (2022); Li et al. (2023), or recent vision language models Li et al. (2023); Liu et al. (2024b); Team (2024a); Bai et al. (2023), all of these methods are built upon large scale high-quality visual understanding data. LAION Schuhmann et al. (2021, 2022), TaiSu Liu et al. (2022), Coyo Byeon et al. (2022), and etc. provide relatively low-quality alt-texts paired with source images. Subsequent works such as CapsFusion Yu et al. (2024), SA1B-Recaption Data (2024b), DataComp-Recaption Li et al. (2024b), and etc. explore enhancing the caption data quality with recaption techniques, and the resulting datasets are widely used in recent VLM researches.
6.2 Vision Language Model Pretrain
VLM pretraining benefits from higher-quality and larger-scale visual understanding data effectively. Previous works, such as BLIP2 Li et al. (2023) and LLaVA Liu et al. (2024b), pretrain the model with relatively low-quality caption datasets Li et al. (2023). With the advance in data engineering works, MiniCPM-V Yao et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c), and QwenVL Bai et al. (2023); Wang et al. (2024) series explore expanding high-quality visual understanding data sizes to improve model performance. In this work, we futher reveal model performance dynamics w.r.t. data quality and training data size, which are largely unexplored in previous works.
6.3 Visual Instruction Tuning
LLaVA Liu et al. (2024b) first defines visual instruction tuning and provides a baseline for VLM SFT data curation. Subsequent LLaVA series models Liu et al. (2023, 2024a); Li et al. (2024a) refine the visual instruction tuning datasets and achieve significantly better model performance. BLIP3 Xue et al. (2024) incorporates image-text interleaved data into visual instruction tuning, while CogVLM Hong et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c) and QwenVL series Wang et al. (2024) models explore using video question answering data for VLM SFT. In this paper, we elaborate the guidelines for the design of visual instruction datasets, providing valuable references for VLM training.
7 Conclusions
In this work, we introduce SAIL-VL-2B, a compact opensource vision language model with SOTA visual comprehension performance. We propose a scalable caption data construction pipeline, and curate SAIL-Caption, a caption dataset with large quantity and the highest quality among opensource alternatives. Pretraining on SAIL-Caption, we demonstrate that even compact VLMs can benefit from scaled up training data size. We further present data size scaling laws that SAIL-VL’s visual comprehension performance improves logarithmically as training data size increases. For visual instruction tuning, we elaborate several key guidelines for SFT data curation, guided by which we curate SFT datasets with balanced distribution. Expected data size scaling laws are observed during SFT on the curated datasets, and the phased SFT strategy further improves the logarithmic scaling curves to be near-linear. We evaluate SAIL-VL-2B on 18 opensource VLM benchmarks, and SAIL-VL outperforms existing VLMs of comparable sizes consistently either in overall performance or domain-specific abilities, depicting promising prospects in real-world applications.
8 Limitations
Despite the leading performance of SAIL-VL among VLMs of comparable sizes, we acknowledge the potential insights that could be gained from experimenting with larger models. We intend to explore this avenue in future work to enhance the robustness of the presented data size scaling laws and other findings. Additionally, our exploration of data size scaling laws has been confined to a specific data magnitude. Although model performance is observed to be saturating at this data quantity, it remains uncertain whether there is room for further improvement under optimized training settings.
Acknowledgments
Our model is built upon many outstanding open-source projects, and we are grateful for their contributions. We extend special thanks to the InternVL team, Qwen team for their great base models, and to the BAAI team (Infinity-MM), Molmo team for their generous release of data.
Appendix A Authorship and Credit Attribution
Data Construction
Hongyuan Dong: scalable data annotation pipeline, caption data construction, and data distribution evaluation
Weijie Yin: data collection and caption data construction
Pretraining
Weijie Yin: Pretrain-alignment stage, data distribution ablation
Hongyuan Dong: Pretrain-advance stage, data distribution ablation, VLM visual understanding ability evaluation, data size scaling laws
SFT
Zijiang Kang: SFT-instruction and SFT-preference, SFT data curation guidelines, phased SFT strategy and SFT data size scaling laws
Weijie Yin: SFT-knowledge and SFT-preference, phased SFT strategy and SFT data size scaling laws
Evaluation
Weijie Yin: data collection, evaluation pipeline
Project Lead
Xiao Liang, Chao Feng and Jiao Ran: resource allocation, project management.
Appendix B SAIL-Caption Showcase
B.1 Caption Data Quality Assessment
To evaluate the caption data quality of SAIL-Caption and opensource alternatives, we curate an evaluation subset for each recaption dataset as the test set for evaluation. Specifically, we randomly select 500 samples from DataComp and SA1B, respectively, and ask GPT4O OpenAI (2024) to generate ground truth captions for subsequent evaluation. SAIL-Captioner is tasked to annotate detailed captions for these samples, and the annotations are evaluated by GPT4O according to the precision and recall of visual elements. We set SA1B QwenVL Caption Data (2024b) and DataComp LLaVA Caption Li et al. (2024b) as baseline datasets, and their data quality is evaluated within the same suite. SA1B QwenVL Caption Data (2024b) employs a finetuned QwenVL model to annotate Chinese dense caption on the SA1B dataset. DataComp LLaVA Caption Li et al. (2024b), on the other hand, trains a customized version of LLaVA model to perform annotation. As shown in Table 6, SAIL-Caption-DataComp and SAIL-Caption-SA1B achieve significantly higher scores than previous baseline datasets, demonstrating its advantage in data quality.
Dataset Language # Sample Captioner Qualtiy Score
[…]
Table 6: Data quality evaluation results of SAIL-Caption and other opensource caption datasets. We curate a 500-sample subset for each dataset for evaluation.
B.2 SAIL-Caption Showcases
Figure 7: SAIL-Caption showcases. Image and questions are curated from SA1B Kirillov et al. (2023) and DataComp Gadre et al. (2024). Highlighted texts denote visual elements depicted by SAIL-Caption but neglected in other caption datasets.
We curate several image samples from SA1B Kirillov et al. (2023) and DataComp Gadre et al. (2024) as demonstrations to compare the quality of SAIL-Caption with existing opensource caption datasets. We compare SAIL-Caption with SA1B-QwenVL-Caption and DataComp-LLaVA-Caption, and showcases are shown in Figure 7. As the demonstrations show, SAIL-Caption encompasses more detailed visual elements than other alternative datasets in both English and Chinese, which coincides with quantified caption quality evaluation results shown in Section 3.1, underscoring its superior data quality.
Appendix C SAIL-VL Showcases
Figure 8: SAIL-VL showcases. Image and questions are curated from LLaVABench Li et al. (2024a).
We task SAIL-VL to tackle vision-based language interactions in Figure 8. The images and input questions are selected from LLaVABench Li et al. (2024a). SAIL-VL demonstrates marked capabilities in language interactions and visual comprehension of various input visual element types, demonstrating powerful visual instruction following abilities.
Appendix D Visual Understanding Benchmark
To inspect SAIL-VL’s visual understanding performance during pretraining stages, we curate a series of visual understanding benchmarks for evaluation. To be specific, we focus on evaluating model performance in detailed captioning and OCR tasks in both English and Chinese, which are also the optimization objectives of pretraining stages. We list the basic information of the selected visual understanding benchmarks in Table 7.
Benchmark Task type Language # Sample
DetailCaps-4870-EN Dong et al. (2024) Caption EN 4870
DetailCaps-4870-CN Dong et al. (2024) Caption CN 4870
IDL-WDS Biten et al. (2022) OCR EN 1000
DocStruct Wang et al. (2020) OCR EN 1000
SynthText Gupta et al. (2016) OCR EN 1000
SynthDog-EN Kim et al. (2022) OCR EN 1000
SynthDog-ZH Kim et al. (2022) OCR CN 1000
Table 7: Basic information of the visual understanding benchmarks used in our experiments.
Benchmarks.
DetailCaps-4870 Dong et al. (2024) encompasses images from a wide range of publicly available datasets, including COYO Byeon et al. (2022), LAION Schuhmann et al. (2021), CC Changpinyo et al. (2021), Flickr Young et al. (2014), SBU Ordonez et al. (2011), and COCO Chen et al. (2015), as well as ground truth detail captions generated by powerful VLM APIs. To ensure evaluation reliability, we task human annotators to refine this dataset by removing hallucinatory descriptions, correcting descriptive errors, and supplementing missing details. The dataset correction is fulfilled in Chinese, with GPT4O OpenAI (2024) deployed by Microsoft Azure as the translator. We adopt both the corrected Chinese captions and translated English captions for evaluation.
The remaining OCR benchmarks consist of images with a diverse distribution. IDL-WDS Biten et al. (2022) consists of document pages with abundant text information; DocStruct Wang et al. (2020) contains both document pages but also illustrative images rendered from tables and charts; SynthText Gupta et al. (2016) is composed of images with a single word, but the fonts vary from one sample to another; SynthDog-EN and SynthDog-ZH Kim et al. (2022) are compositional datasets comprised of natural image backgrounds and foreground word pieces.
Metrics.
We evaluate SAIL-VL’s caption performance on the DetailCaps-4870 benchmark with GPT evaluation. Provided with three ground truth captions and a candidate caption, GPT is tasked to score the candidate caption based on the precision and recall of the visual elements. For OCR tasks, we compute the ANLS score Biten et al. (2019) between the predicted OCR contents and the ground truth one, resulting in scores ranging from 0 to 1, with higher score indicating better prediction quality.
All benchmark data is curated from left-out subsets to avoid data leakage between model training and evaluation. We select a 500-case subset randomly from each benchmark to evaluate the pretrained model checkpoints for efficiency.